a 2004 Pontiac Montana has an engine coolant system capacity of 10.01 liter. If the system is currently filled with a mixture that is 30% antifreeze, how much of it should be drained and replaced with pure antifreeze so that they system is filled with a mixture that is 50% antifreeze?
Answer by stanbon(75887)

Pontiac Montana has an engine coolant system capacity of 10.01 liter. If the system is currently filled with a mixture that is 30% antifreeze, how much of it should be drained and replaced with pure antifreeze so that they system is filled with a mixture that is 50% antifreeze?
-----------------------
Let "x" be the amount that should be drained and replaced.
------------------------------
Equation:
0.30*10.01 - 0.30x + 1.00x = 0.50*10.01
---
Multiply thru by 100 to get:
30*10.01 + 70x = 50*10.01
70x = 20*10.1
x = 2.89 liters (amount that should be drained and replaced)
================================================================
Cheers,
Stan H.